OpenStudy (anonymous):

Solve -5w = -80 for w.

OpenStudy (anonymous):

16

OpenStudy (anonymous):

Solve d - (-2) = 19 for d.

OpenStudy (anonymous):

17

OpenStudy (anonymous):

Solve -12 = z ÷ 3 for z.

OpenStudy (anonymous):

z=

OpenStudy (shaik0124):

w=80/5 w=16

OpenStudy (anonymous):

-24

OpenStudy (anonymous):

-24/3 = -12

OpenStudy (anonymous):

what one is -12

OpenStudy (anonymous):

Solve -12 = z ÷ 3 for z. z=

OpenStudy (shaik0124):

-12*3=-36 is answer